OpenStudy (anonymous):

Solve for y to obtain the slope-intercept form: 5x + y =12 y= ? x - y = 12 y= ?

OpenStudy (anonymous):

can you please specify exactly what you want?

OpenStudy (anonymous):

... The slope-intercept form of both equations, well I would like to know how to go about finding the slope-intercept form.

OpenStudy (anonymous):

y = -5x +12 and y = x-12 This is the slope intercept form y=mx+c where m is the slope and c is the intercept.

OpenStudy (anonymous):

OH okay! I see. Thank you.